OpenStudy (itskayylayy1213):

x + y = −3 y = 2x + 2 HELPPPPP

OpenStudy (rebecca):

Well we know that y=2x+2 therefore plug that in for the first problem x+(2x+2)=-3

OpenStudy (rebecca):

Now solve for x.

OpenStudy (itskayylayy1213):

i honestly have no idea hoe to do this . like at all

OpenStudy (itskayylayy1213):

how*

OpenStudy (rebecca):

Add the x terms. Then move all constants to the other side.

OpenStudy (itskayylayy1213):

yeah i have no idea what constants are

OpenStudy (rebecca):

Example 2x+(3x+1)=-4 2x+3x=-4-1 5x=-5 x=-1

OpenStudy (itskayylayy1213):

can you help me with this one -_- please

OpenStudy (rebecca):

I'm giving you an example. I can't do the work for you.

OpenStudy (itskayylayy1213):

im aware i mean guide me ...

OpenStudy (rebecca):

Okay. What is x+2x=?

ILovePuppiesLol (ilovepuppieslol):

she's guiding you ^_^ don't worry, just be patient!

OpenStudy (rebecca):

1x=x so 1x+2x=?

OpenStudy (itskayylayy1213):

i dont know nvm ill guess

ILovePuppiesLol (ilovepuppieslol):

aw guessing doesnt help u learn :(

OpenStudy (rebecca):

What is 1+2=?

ILovePuppiesLol (ilovepuppieslol):

21

OpenStudy (xxdragoonxx):

x + y = −3 x+y+−y=−3+−y x=−y−3 y=2x+2 2x+2=y 2x+2+−2=y+−2 2x=y−2 2x/2 = y-2/2 x = 1/2y-1